OpenStudy (anonymous):

Please Solve Simultaneous Equasion 2y+x=4 3y-2x=13

OpenStudy (anonymous):

do i have to show u ?

OpenStudy (anonymous):

In the first equation, isolate x. Then, plug that value into the x value of the second equation to find y. With the y value, you can find x by plugging it into either of the equations assuming that this is a system of equations.

OpenStudy (anonymous):

2y + x=4 x = 4 - 2y 1st eq then replace the value in second equation 3y - 8 + 4y =13 7y=21 y=3 then again replace this valuein 1st eq x = 4 -6 x =-2

OpenStudy (amistre64):

2y + x = 4 ; multiply by 2 3y -2x = 13 4y +2x = 8 3y -2x = 13 ; now add together ------------ 7y + 0 = 21 7y = 21 ; solve for y and use it to solve for x

OpenStudy (anonymous):

interesting tactic amistre by multiplying by 2 at the start!

OpenStudy (anonymous):

Ishaaan, you didn't need to show me, I am just confirming my answers...cheers!

OpenStudy (amistre64):

thanx, its a method called elimination of a variable.

OpenStudy (anonymous):

can i ask, why multiply by 2 amistre?